In a regression analysis involving 30 observations, the following estimated regression equation was obtained.

ŷ = 16.6 + 3.5x1 − 2.2x2 + 7.8x3 + 2.8x4

a) Predict y when x1 = 10, x2 = 5, x3 = 1, and x4 = 2.

ŷ = 17.9 + 3.6x1 − 2.2x2 + 7.9x3 + 2.8x4

b)Predict y when x1 = 10, x2 = 5, x3 = 1, and x4 = 2.
Answer #1

(Date: 1 6) Here y = 16.6*3.5*, - 2.282+7.8 X3 +2.8X4 Fer x1=10, X2=5 X = 1 , Xa=2 y = 16.6 +3.5X10 - 212X5+ 7.8*l+2.872 6) ИDear student,
I am waiting for your feedback. I have given my 100% to solve your queries. If you are satisfied by my given answer. Can you please like it☺